A cannon ball is projected so as to attain a maximum range. Find the maximum height attained if the initial velocity is u
1
Expert's answer
2020-04-27T09:59:29-0400

A cannon ball is projected so as to attain a maximum range. Thus, the angle is


"\alpha=45\degree"

So


"H=\frac{u^2\sin^2{45}}{2g}=\frac{u^2(0.5)}{2g}"

"H=\frac{u^2}{4g}"